# Exploring run-and-tumble movement in confined settings through simulation

### v1

@@ 4 Distributions of path length between collisions
Figure 8: (a) Variance of the path length versus the number of collisions with elastic reflecting conditions at the boundary. (b) Variance of the path length versus the parameter $l/R$ for simulations with $10^{9}$ collisions with elastic and random reflecting conditions at the boundary in log-log representation.
@@
Hereinafter, we will focus in the particular case of exponential step length distribution and uniformly random angular distribution. In the subsequent analysis, we will further explore the dynamic properties of the path length distributions by examining the variance of these distributions. The variance provides crucial insight into the spread of the path length around its mean, which in this context is constant across all distributions at $<L>=\frac{\pi}{2}$ for the elastic boundary conditions, and variable as showed in Fig. (7) for the random angle boundary condition. In Fig. (8), we plot the variance of the path length as a function of two parameters: the number of collisions for the elastic condition, and the parameter $l/R$ for both, elastic and random boundary conditions. This plot illustrates how the variance evolves as the motion transitions from diffusive to quasi-ballistic regimes across different scaling of $l/R$. The limit as $l/R\rightarrow\infty$ is $Var(L)_{elastic}=8/3-\pi^{2}/4\sim 0.199$ for the elastic case. On the contrary, the random billiard has the limit $Var(L)_{random}=2-16/\pi^{2}\sim 0.378$. These limits can be obtained by a straightforward calculation. For the elastic case, the measure is $\mu(d\theta)=\frac{1}{2}\cos\theta d\theta$, and, by the geometry of the problem, a chord between two points in the boundary has length $L=2R\cos\theta$, see Fig. (14(a)), where $\theta$ is the outgoing angle from the first point to the second one, measured with respect to the normal. Therefore,
$<L^{2}>_{elastic}=\int_{-\pi/2}^{\pi/2}(2R\cos\theta)^{2}\times\frac{1}{2}\cos\theta\times d\theta=\frac{8}{3}R^{2},$ (6)
@@
and the variance for our case $R=1$ is $Var(L)_{elastic}=<L^{2}>-<L>^{2}=8/3-\pi^{2}/4$. On the other hand, in the case of random boundary condition, the measure is $\mu(d\theta)=d\theta/\pi$, corresponding to a collision rule $P_{R}(\theta)=1/\pi$, the first moment is $<L>_{random}=4/\pi$ [32], and the second moment is
@@
Figure 9: Path length distribution with elastic boundaries, $10^{7}$ collisions, 100 bins. (a) $l/R=0.01$, (b) $l/R=0.1$, (c) $l/R=1$, (d) $l/R=10$. It can be seen how the distribution evolves from diffusive (a) to quasi-ballistic (d).
A more clear picture can be made by looking at the distributions. In Fig. (9), the distribution of path length between collisions is shown, featuring elastic boundaries, and the conditions previously mentioned, for various values of the parameter $l/R$: (a) $l/R=0.01$, (b) $l/R=0.1$, (c) $l/R=1$, (d) $l/R=10$. As illustrated in Fig. (2), while all distributions share the same mean, $<L>=\pi/2$, one can observe how the distribution transitions from diffusive (a) to quasi-ballistic (d).
Figure 10: Path length distribution with elastic boundaries, $10^{7}$ collisions, 100 bins. Ratio $l/R=0.1$. (a) linear, (b) semilog, and (c) log-log representation. Red lines are exponential fittings, green lines are power law fittings. It can be appreciated that the distribution is a combination of both functions.
Interestingly, when $l/R<1$, the distribution is an exponential decay with a power law deviation for small values of length, as illustrated in Fig. (10) and (11). The distributions are presented in (a) linear, (b) semilog, and (c) log-log formats. It is noteworthy that for $l/R=0.3$, the semilog representation displays a straight line, whereas for $l/R=0.05$, it is the log-log representation that appears straight. If we decrease the ratio $l/R$, the power law deviation for small lengths grows.
@@
Figure 12: Cumulative probability distributions of path length between collisions for (a) $l/R=0.002$ and (b) $l/R=1$. Both with elastic boundary conditions, exponential step length, uniformly random angular distribution, $10^{8}$ collisions, and 600 bins.
This is better appreciable by looking at the cumulative distributions. We show the cumulative probability distribution in Fig. (12) for two different ratios $l/R$. The distributions were made with $10^{8}$ collisions and 600 bins. It is evident that, excluding small lengths, an exponential decay, $P=P_{0}\exp(-\gamma\ell)$, can be fitted. The parameter of the decay, $\gamma$, depends on $l/R$, as shown in Fig.(13). A parabolic fit seems to represent the numerical data quite well, but we have no theoretical argument to support this choice.
Figure 13: Exponent $\gamma$ of the exponential fit of cumulative distributions as a function of $l/R$. The red line is a parabolic fit of the points.
When $l/R>1$, the distributions, Fig. (9(d)) for example, are a combination of a diffusive and a ballistic distribution, since sometimes the step length is equal or greater than $R$, contributing to the ballistic mode, and other times the step length is small and therefore contributes to the diffusive mode. Let us name $\tau$ the length of a chord between two points of the boundary. Remembering that $\tau=2R\cos\theta$, see Fig. (14(a)), where $\theta$ is the outgoing angle from the first point to the second one, measured with respect to the normal. Then we have
@@
As can be seen in Fig. (14(b)), the ballistic part can be descripted with high accuracy by eq. (10). In this figure, we show the ballistic part of the path length distribution for elastic boundary conditions, exponential step length distribution and uniformly random angular distribution. The distribution is constructed with 100 bins and $10^{7}$ collisions, for the case $l/R=10$.
@@ 5 Radial distributions
We performed the simulations with exponential step distribution and uniformly random angular distribution. The radial distribution was found to be linear under elastic boundary conditions for all values of the ratiof $l/R$. For instance, in Fig. (15(a)), we show the distribution for the case $l/R=0.3$. This can be analytically calculated as follows. Imagine a chord between points at the boundary representing a part of the trayectory of the particle, as represented in Fig. (14(a)). The probability of being at a given point on the chord, $z$, at a distance $r$ from the center is
@@
where $\theta$ is the angle measured with respect to the normal, see Fig. (14(a)). By geometry we get
@@
$<\tau>=R\int_{-\pi/2}^{\pi/2}cos^{2}\theta.d\theta=\frac{\pi R}{2},$ (15)
and therefore, the radial probability can be calculated as
$p_{elastic}(r)=\int\frac{\tau}{<\tau>}W_{\theta}(r)\cos\theta.d\theta=\frac{4r}{\pi R}\int_{0}^{arcsin(r/R)}\frac{\cos\theta.d\theta}{\sqrt{r^{2}-R^{2}sin^{2}\theta}}$ (16)
@@
$\cos\phi.d\phi=\frac{R}{r}\cos\theta.d\theta$ (18)
@@
$p_{elastic}(r)=\frac{4r}{\pi R}\int_{0}^{\pi/2}\frac{\cos\phi.d\phi}{r\cos\phi}=\frac{2r}{R^{2}}.$ (19)
@@
In contrast, for random angle boundary conditions, the radial distribution deviates from linearity and is described by the complete elliptic integral of the first kind that can be calculated as follows. In the case of random angle boundary conditions, the measure is $d\theta/\pi$, and therefore,
$<\tau>=\frac{2R}{\pi}\int_{-\pi/2}^{\pi/2}cos\theta.d\theta=\frac{4R}{\pi}.$ (21)
@@
$p_{random}(r)=\int\frac{\tau}{<\tau>}W_{\theta}(r)\frac{2}{\pi}.d\theta=\frac{r}{R}\int_{0}^{arcsin(r/R)}\frac{d\theta}{\sqrt{r^{2}-R^{2}\sin^{2}\theta}}.$ (22)
@@
$p_{random}(r)=\frac{r}{R^{2}}\int_{0}^{\pi/2}\frac{d\theta}{\sqrt{1-\frac{r^{2}}{R^{2}}sin^{2}\theta}}=\frac{r}{R^{2}}K\left(\frac{r}{R}\right),$ (23)
where $K(x)$ is the complete elliptic integral of the first kind, defined as [39]
$K(x)=\int_{0}^{\pi/2}\frac{ds}{\sqrt{1-x^{2}sin^{2}s}}.$ (24)
This expression is compared with the simulation in Fig. (15(b)) for the case $l/R=10$ and, as can be seen, the agreement is satisfactory. An alternative procedure for calculating these distributions are presented in the appendix. This distribution shows that particles are more likely to be found near the walls of the domain under random angle reflections, these results seem to be in agreement with experiments with microbia in which the microbe spends the majority of their time close to the walls of the domain [37, 40]. In fact, the study of the radial distribution can bring information about the mobility mechanism in experimental settings, since different distributions are generated depending on how the active agent interacts with the walls and, as we will show in the next section, the reorientation mechanism. It would be interesting, therefore, to see more experimental outcomes on this matter. A related concept, the boundary local time, i.e. how long a diffusing particle spends in a close vicinity of a confining boundary, plays thus a crucial role in the description of various surface-mediated phenomena, such as heterogeneous catalysis, permeation through semipermeable membranes, or surface relaxation in nuclear magnetic resonance [41].
6 Processes with memory
Let us now focus in a different, non-trivial case. Process memory refers to the phenomenon where the future state of a particle’s motion is influenced by its past states, leading to non-Markovian dynamics. The concept of process memory plays a crucial role in understanding the reorientation mechanism of active agents in run-and-tumble movement. For example, some processes like chemotaxis, introduce a kind of memory, namely, the ability of the particle to compare the concentration or the gradient of a chemical in different steps of its movements.

### v2

@@ III Mean path length theorem
Figure 6: (a) Mean path length versus number of collisions with elastic boundaries. Each point is an average of 20 simulations. (b) Mean path length versus the number of collisions with random reflecting conditions at the boundary.
Let’s now examine the scenario where the outgoing angle after a collision with the boundary is random. Initially, the absolute values of the mean path length under elastic boundary conditions, with an exponential step length distribution and uniformly random angle distribution, are presented in Fig. (6(a)). These values approach the theoretical value of $\frac{\pi}{2}=1.57079...$ as the number of collisions approaches infinity. For contrast, the absolute values of the mean path length for the random reflecting angle condition are plotted in Fig. (6(b)). As can be seen, the invariant property no longer holds, as we anticipated in Artuso and Zamora (2024). On the contrary, the mean path length does depend on $l/R$, i.e., the type of movement (diffusive or ballistic), as can be observed in Fig. (6(b)).
Figure 7: Mean path length as a function of $l/R$, $10^{9}$ collisions, random angle boundary conditions. The limit for $l/R\rightarrow\infty$ is the mean path length of a random billiard, $1.27324/\pi=$ 1.2732 …, represented by the black line.
In Fig. (7), we show this dependency in a $<L>$ versus $l/R$ plot for simulations with $10^{9}$ collisions. It can be seen that the mean path length tends to $4/\pi$, which is the mean path length of a random billiard as shown in Artuso and Zamora (2024). This behavior is consistent with the underlying physics of the system. In the limit where $l/R$ becomes very large—referred to as the ballistic regime—the particle’s motion becomes dominated by long, straight trajectories between collisions with the boundary. In this regime, the particle travels in nearly straight lines across the domain, and its motion closely resembles that of a random billiard, where the particle’s path is only altered by reflections at the boundary. For a random billiard, it has been shown that the mean path length, $<L>$, is exactly $4/\pi$. This convergence reflects the transition from diffusive behavior, characterized by shorter and more varied paths, to ballistic behavior, where the mean path length stabilizes at the expected value for a random billiard.
The MPLT holds true across various step length distributions—whether Gaussian, uniformly random, fixed, or exponentially distributed—as these do not affect the uniformity of particle incidence on the boundary, a critical assumption of the theorem. However, altering the angular distribution of the tumble events or the angle post-collision breaks this assumption. When the angular distribution deviates from being uniformly random, it introduces anisotropy or non-uniformity in the way particles interact with the boundary, violating the MPLT’s conditions. For instance, fixed or Gaussian-distributed angles after tumble events create preferred directions of movement, leading to non-homogeneous wall collisions and thereby causing the theorem to fail. The MPLT does not hold under random boundary conditions, as opposed to elastic conditions, due to the way these conditions alter the uniformity and isotropy of particle interactions with the boundary of the domain. In contrast, under elastic boundary conditions, a particle’s angle of reflection equals its angle of incidence, preserving the distribution of angles at which particles strike and leave the boundary. This uniformity ensures that the particle flux across the boundary maintains the assumptions required for the MPLT to apply.
IV Distributions of path length between collisions
In the context of boundary interactions, homogeneous flux of particles at the boundary refers to a situation where particles strike the boundary uniformly, meaning that every point on the boundary has an equal likelihood of being hit by a particle. Isotropic flux means that particles approach the boundary from all directions with equal probability, ensuring that there is no preferred direction in the incidence of particles. After the evidence we showed in the last section, it is evident that under the condition of homogeneous and isotropic flux of particles at the boundary, the mean path length is determined by Cauchy’s formula. However, this prompts the question: how does the distribution behave under different conditions? Let us delve into this inquiry.
Figure 8: (a) Variance of the path length versus the number of collisions with elastic reflecting conditions at the boundary. (b) Variance of the path length versus $l/R$, $10^{9}$ collisions. The dashed lines are the variances of elastic and random billiards.
@@ IV Distributions of path length between collisions
Hereinafter, we will focus on the particular case of exponential step length distribution and uniformly random angular distribution. In the subsequent analysis, we will further explore the dynamic properties of the path length distributions by examining the variance of these distributions. The variance provides crucial insight into the spread of the path length around its mean, which in this context is constant across all distributions at $<L>=\frac{\pi}{2}$ for the elastic boundary conditions, and variable as shown in Fig. (7) for the random angle boundary condition. In Fig. (8), we plot the variance of the path length as a function of two parameters: the number of collisions for the elastic condition, and the parameter $l/R$ for both, elastic and random boundary conditions. This plot illustrates how the variance evolves as the motion transitions from diffusive to quasi-ballistic regimes across different scalings of $l/R$. The limit as $l/R\rightarrow\infty$ is $Var(L)_{elastic}=8/3-\pi^{2}/4\sim 0.199$ for the elastic case. On the contrary, the random billiard has the limit $Var(L)_{random}=2-16/\pi^{2}\sim 0.378$. These limits can be obtained by the following calculation.
Let first consider usual billiards. Usual billiards involve elastic (specular) reflections: the ingoing and outgoing angles coincide, or, equivalently
$v_{1}=v_{0}-2(n(\tilde{q})\cdot v_{0})n(\tilde{q}),$ (6)
where $v_{0}$ and $v_{1}$ are the incoming and outgoing velocities, respectively, $\tilde{q}$ is the point of impact along the boundary, and $n(\tilde{q})$ is the (inward) normal to the boundary at the point $\tilde{q}$. Notice that elastic reflection, besides being dictated by elastic collision law with an infinite massive boundary segment, makes the billiard flow a Hamiltonian system Chernov and Lebowitz (1997); Bonetto, Chernov, and Lebowitz (1998): in particular if $\phi$ denotes the ingoing angle, and $\psi=f(\phi)$ the outgoing angle (for a deterministic collision rule) there is a phase space contraction $\alpha$ at each collision event, with
$\alpha=-\log\left(\frac{\sin\psi}{\sin\phi}f^{\prime}(\phi)\right).$ (7)
So, specular reflections make the billiard flow a Hamiltonian system, for which volume in the phase space is conserved. Even if the following considerations are valid in any dimension, we will consider billiard tables in a bounded region $\Omega$ of $\mathbb{R}^{2}$: due to conservation of kinetic energy the phase space is thus $\Gamma=S^{1}\times\Omega$, and the (invariant) Lebesgue measure is
$d\mu=C_{\mu}\cdot d\theta\,dx\,dy\qquad C_{\mu}=(2\pi V_{\Omega})^{-1},$ (8)
where $(x,y)$ denote spatial coordinates in $\Omega$, and $\theta$ is the angle denoting the direction. This invariant measure defines, for any flow observable, a phase average, which coincides a.e. with time averages provided the flow is ergodic (incidentally this is far from being trivial for large classes of convex billiards). The connection to billiard maps (discrete dynamical system with phase space coordinates $\tilde{q},\phi$), is made clear by a change of coordinates in the space phase volume: if we fix a point $x,y,\theta$, then the backward flow uniquely determines $\tilde{q},\phi$ (coordinates for the billiard map), together with $\tau_{-}(x,y,\theta)$ (time from the boundary to the point $x,y,\theta$): so if we change $dx,dy$ into the product of a line element along the flow ($dt$) and another one perpendicular to it ($\cos(\phi)\,d\tilde{q}$) Sinai (1970); Chernov (1991, 1977), we get
$d\mu=C_{\mu}\cdot d\phi\,\cos(\phi)\,d\tilde{q}\,dt,$ (9)
where $d\tilde{q}$ is the Lebesque measure along the boundary $\partial\Omega$. This is associated to an invariant measure for the billiard map:
$d\nu=C_{\nu}\cdot d\tilde{q}\,\cos(\phi)d\phi\qquad C_{\nu}=(2\cdot S_{\partial\Omega})^{-1}.$ (10)
Moreover we can easily get Cauchy formula for billiards:
[…]
namely, the phase average of the collision to collision time (averaged with the invariant measure for the billiard map), obeys (in dimension $d=2$)
$\langle\tau\rangle_{\nu}=\frac{C_{\nu}}{C_{\mu}}=\pi\frac{V_{\Omega}}{S_{\partial\Omega}}.$ (12)
This identity establishes Cauchy law (for the average of collision to collision times) for ergodic billiards (irrespective of geometric properties of the shape of the billiard table). While in former considerations the cosine law (10) arises as the invariant measure of a fully deterministic system, it was assumed as the distribution of reflected gas molecules irrespective of their incident angle by Knudsen Knudsen (1952). Knudsen law has been discussed as arising from microstructure of the billiard boundary in Feres and Yablonsky (2004), while Celestini and Mortessagne (2008) investigate how it statistically arises from microscopic molecular simulations.
After these considerations, replacing the value of $C_{\mu}$ for the particular case of a circle in the cosine law, we obtain that, for the elastic case, the measure is $\mu(d\theta)=\frac{1}{2}\cos\theta d\theta$, and, by the geometry of the problem, a chord between two points on the boundary has length $L=2R\cos\theta$, see Fig. (13(a)), where $\theta$ is the outgoing angle from the first point to the second one, measured with respect to the normal. Therefore,
$\begin{split}<L^{2}>_{elastic}&=\int_{-\pi/2}^{\pi/2}(2R\cos\theta)^{2}\times\frac{1}{2}\cos\theta\times d\theta\\ &=\frac{8}{3}R^{2},\end{split}$ (13)
@@
and the variance for our case $R=1$ is $Var(L)_{elastic}=<L^{2}>-<L>^{2}=8/3-\pi^{2}/4$. On the other hand, in the case of random boundary conditions, the measure is $\mu(d\theta)=d\theta/\pi$, corresponding to a collision rule $P_{R}(\theta)=1/\pi$. The first moment is $<L>_{random}=4/\pi$ Artuso and Zamora (2024), and the second moment is
@@
Figure 9: Path length distribution with elastic boundaries, $10^{7}$ collisions, 100 bins. (a) $l/R=0.01$, (b) $l/R=0.1$, (c) $l/R=1$, (d) $l/R=10$.
A clearer picture can be obtained by examining the distributions. In Fig. (9), the distribution of path lengths between collisions is shown, featuring elastic boundaries and the conditions previously mentioned, for various values of the parameter $l/R$: (a) $l/R=0.01$, (b) $l/R=0.1$, (c) $l/R=1$, (d) $l/R=10$. As illustrated in Fig. (2), while all distributions share the same mean, $<L>=\pi/2$, one can observe how the distribution transitions from diffusive (a) to quasi-ballistic (d).
Figure 10: Path length distribution with elastic boundaries, $10^{7}$ collisions, 100 bins. Ratio $l/R=0.1$. (a) linear, (b) semilog, and (c) log-log representation. Red lines are exponential fittings, green lines are power-law fittings.
@@
Interestingly, when $l/R<1$, the distribution exhibits an exponential decay with a power-law deviation for small values of length, as illustrated in Figs. (10) and (11). The distributions are presented in (a) linear, (b) semilog, and (c) log-log formats. It is noteworthy that for $l/R=0.3$, the semilog representation displays a straight line, whereas for $l/R=0.05$, it is the log-log representation that appears straight. As we decrease the ratio $l/R$, the power-law deviation for small lengths becomes more pronounced. The decay parameter of the exponential fitting, $\gamma$, depends on $l/R$, as shown in Fig. (12). A quadratic fitting seems to represent the numerical data quite well, but we have no theoretical argument to support this choice. The best values of the fit are presented in the label of Fig. (12).
Figure 12: Decay parameter $\gamma$ of the exponential fitting as a function of $l/R$. The red line is a quadratic fit of the points.
When $l/R>1$, the distributions, as shown in Fig. (9(d)), for example, are a combination of diffusive and ballistic distributions. This occurs because sometimes the step length is equal to or greater than R, contributing to the ballistic mode, while at other times the step length is small and therefore contributes to the diffusive mode. Let us denote by $\tau$ the length of a chord between two points on the boundary. Recall that $\tau=2R\cos\theta$, see Fig. (13(a)), where $\theta$ is the outgoing angle from the first point to the second, measured with respect to the normal. Then we have
@@
As can be seen in Fig. (13(b)), the ballistic part can be described with high accuracy by Eq. (17). In this figure, we show the ballistic part of the path length distribution for elastic boundary conditions, an exponential step length distribution, and a uniformly random angular distribution. The distribution is constructed with 100 bins and $10^{7}$ collisions, for the case $l/R=10$.
@@ V Radial distributions
We performed the simulations with an exponential step length distribution and a uniformly random angular distribution. The radial distribution was found to be linear under elastic boundary conditions for all values of the ratio $l/R$. For instance, in Fig. (14(a)), we show the distribution for the case $l/R=0.3$. This can be analytically calculated as follows. Imagine a chord between points at the boundary representing part of the trajectory of the particle. This is represented by the segment AB in Fig. (13(a)). The probability, $W$, of being between a given point on the chord, $z$, and $z+dz$, is equal to the ratio between $dz$ and the length of the whole chord, $\tau$.
$W_{\theta}(r)dr=\frac{dz(r)}{\tau},$ (18)
where $\theta$ is the angle measured with respect to the normal at the point A. Looking at the triangle ABC and taking into account that the segment AC is the diameter of the circle, we get that the length of the chord AB is $\tau=2R\cos\theta$. Therefore,
@@
where $r$ is the distance from the center of the circle to $z$, see Fig. (13(a)). By geometry of the interior triangle we get
@@
[…]
Suppose you reconstruct the radial density by using the dynamics with discretized steps: then short segments contribute less points than long segments: this can be taken into account exactly by a $\frac{\tau}{<\tau>}$ factor, and therefore, the radial probability can be calculated as
$\begin{split}p_{elastic}(r)&=\int\frac{\tau}{<\tau>}W_{\theta}(r)\cos\theta d\theta\\ &=\frac{4r}{\pi R}\int_{0}^{\arcsin(r/R)}\frac{\cos\theta d\theta}{\sqrt{r^{2}-R^{2}\sin^{2}\theta}}.\end{split}$ (24)
@@
$\cos\phi d\phi=\frac{R}{r}\cos\theta d\theta$ (26)
@@
$p_{elastic}(r)=\frac{4r}{\pi R}\int_{0}^{\pi/2}\frac{\cos\phi d\phi}{r\cos\phi}=\frac{2r}{R^{2}}.$ (27)
@@
In contrast, for random angle boundary conditions, the radial distribution deviates from linearity and is described by the complete elliptic integral of the first kind, which can be calculated as follows. In the case of random angle boundary conditions, the measure is $d\theta/\pi$, and therefore,
$<\tau>=\frac{2R}{\pi}\int_{-\pi/2}^{\pi/2}\cos\theta d\theta=\frac{4R}{\pi}.$ (29)
@@
[…]
@@
$p_{random}(r)=\frac{r}{R^{2}}\int_{0}^{\pi/2}\frac{d\theta}{\sqrt{1-\frac{r^{2}}{R^{2}}\sin^{2}\theta}}=\frac{r}{R^{2}}K\left(\frac{r}{R}\right),$ (31)
where $K(x)$ is the complete elliptic integral of the first kind, defined as Gradshteyn and Ryzhik (2007)
$K(x)=\int_{0}^{\pi/2}\frac{ds}{\sqrt{1-x^{2}\sin^{2}s}}.$ (32)
This expression is compared with the simulation in Fig. (14(b)) for the case $l/R=10$, and as can be seen, the agreement is satisfactory. An alternative procedure for calculating these distributions is presented in the appendix. This distribution shows that particles are more likely to be found near the walls of the domain under random angle reflections. These results seem to be in agreement with experiments with microbes, in which the microbes spend the majority of their time close to the walls of the domain Souzy et al. (2022); Ostapenko et al. (2018). However, the experimental observation might have other physical explanations, such as hydrodynamical interactions, or other kind of microbe-wall interactions. The assumption that the interaction of a swimming microbe with a wall may be described as a collision could be subject to criticism and more realistic modelization of the interactions could be introduced. In fact, the study of the radial distribution can provide information about the mobility mechanism in experimental settings, since different distributions are generated depending on how the active agent interacts with the walls and, as we will show in the next section, the reorientation mechanism. It would be interesting, therefore, to see more experimental outcomes on this matter. A related concept, the boundary local time—i.e., how long a diffusing particle spends in the close vicinity of a confining boundary—thus plays a crucial role in the description of various surface-mediated phenomena, such as heterogeneous catalysis, permeation through semipermeable membranes, or surface relaxation in nuclear magnetic resonance Grebenkov (2019).
VI Processes with memory
Let us now focus on a different, non-trivial case. Process memory refers to the phenomenon where the future state of a particle’s motion is influenced by its past states, leading to non-Markovian dynamics. The concept of process memory plays a crucial role in understanding the reorientation mechanism of active agents in run-and-tumble movement. For example, some processes, like chemotaxis, introduce a kind of memory—namely, the ability of the particle to compare the concentration or gradient of a chemical across different steps of its movement.